On excercise 30A, there are some questions I was confused about, the numbers were, 24, 26, 28 and 30. What we were supposed to do was to find the formula, then find the 10th, or 15th or 20th term. Can you help me? This is how the questions come out in the book. The question below is number 24.
24. 1st:5 2nd:14 3rd:27 4th:44 5th:65 State the 10th and 20th term.

Answer by stanbon(48568)

1st:5 2nd:14 3rd:27 4th:44 5th:65 State the 10th and 20th term.
----------------
Taking successive differences you get:
14-5 = 9 ; 27-14 = 13; 44-27 = 17; 65-44 = 21
----------------
Taking successive differences again you get:
13-9 = 4; 17-13 = 4; 21-17=4
----------------
Since the 2nd differences are a constant number, the pattern is quadratic.
Using the points (1,5)(2,14)(3,27)(4,44)(5,65)
you can find the quadratic pattern is y = 2x^2+3x
----------
Then you can find f(10) = 230
and f(20) = 860
======================
Cheers,
Stan H.